# T3: Multi-modal Tailless Triple-Flapping-Wing Robot for Efficient Aerial and Terrestrial Locomotion

### v1

@@ Abstract
Flapping-wing robots offer great versatility; however, achieving efficient multi-modal locomotion remains challenging. This paper presents the design, modeling, and experimentation of T3, a novel tailless flapping-wing robot with three pairs of independently actuated wings. Inspired by juvenile water striders, T3 incorporates bio-inspired elastic passive legs that effectively transmit vibrations generated during wing flapping, enabling ground movement without additional motors. This novel mechanism facilitates efficient multi-modal locomotion while minimizing actuator usage, reducing complexity, and enhancing performance. An SE(3)-based controller ensures precise trajectory tracking and seamless mode transition.
To validate T3’s effectiveness, we developed a fully functional prototype and conducted targeted modeling, real-world experiments, and benchmark comparisons. The results demonstrate the robot’s and controller’s outstanding performance, underscoring the potential of multi-modal flapping-wing technologies for future aerial-ground robotic applications.
@@ I Introduction
The study of flapping-wing robots has garnered significant attention due to their promising applications in areas such as disaster rescue, environmental monitoring, urban exploration, and entertainment performances [1] [2] [3] [4]. However, these robots often face limitations in maneuverability and adaptability in complex environments [5] [6] [7]. To address these challenges, researchers have developed multi-modal flapping-wing robots capable of both aerial and terrestrial locomotion. This advancement enables robots to t diverse terrains and perform a wider range of tasks, thereby enhancing their versatility and effectiveness in real-world applications.
Figure 1: Multi-modal tailless triple-flapping-wing robot T3.
Integrating mechanisms from ground robots has become a common approach to achieving terrestrial mobility for multi-modal flapping-wing robots, due to their well-established reliability. However, equipping these robots with wheels [8] [9] and walking mechanisms [7] [9] [10] [11] often requires additional actuators compared to their flight-only configurations. This not only complicates the control system but also reduces the robot’s reliability and increases manufacturing and maintenance costs [12]. Moreover, this approach often results in inefficient multi-modal locomotion and switching.
Regarding aerial mobility, there is a growing trend in biomimetic research to achieve high levels of mimicry of natural organisms [5] [6]. For instance, some designs replicate insects by employing two sets of flapping mechanisms and two servos, while others mimic insects with two pairs of independent wings and two rotors [8]. Additionally, various bio-inspired designs have been developed, drawing inspiration from larger birds [10], hummingbirds [13], flying squirrels [14], ladybird beetles [15], and others. To achieve high levels of biomimicry and full attitude control in the air, existing studies often involve complex designs with numerous actuators, which can lead to inefficiency in multi-modal locomotion.
@@
Building upon previous analysis, we present a multi-modal tailless flapping-wing robot equipped with bio-inspired elastic passive legs modeled after those of juvenile water striders, which are used for locomotion on the water’s surface [16] [17] [18]. This design leverages the robot’s wing vibrations to enable terrestrial locomotion, integrating various functionalities such as vertical takeoff, multi-degree-of-freedom flight, self-righting, terrestrial locomotion, and seamless mode transition—all achieved with just three actuators. The robot features three sets of annularly symmetric flapping-wing actuators for propulsion, with a total weight of $37.4\,\mathrm{g}$. In flight mode, it reaches a maximum speed of $5.5\,\mathrm{m/s}$, with a flight endurance of $6.5\,\mathrm{minutes}$. In terrestrial mode, it achieves a top speed of $5.4\,\mathrm{cm/s}$ and can sustain controlled locomotion for up to $28\,\mathrm{minutes}$.
Figure 2: (a) Illustration of the terrestrial locomotion modeled after those of juvenile water striders during their locomotion on the water’s surface [16] [17] [19], (b) Bio-inspired elastic passive leg, (c) Water strider’s leg during their locomotion on the water’s surface (typical side view) [18].
@@
This design effectively addresses key challenges in both terrestrial and aerial mobility. By incorporating bio-inspired elastic passive legs, our approach utilizes wing vibrations for ground movement, eliminating the need for additional actuators. This not only simplifies the control system but also enhances reliability while reducing manufacturing and maintenance costs. In terms of aerial mobility, the three-winged configuration optimizes thrust generation and stability while maintaining a lightweight structure. Unlike conventional designs, which rely on numerous actuators for high-performance attitude control, our method reduces system complexity without compromising maneuverability. This innovative approach improves the efficiency of multi-modal locomotion and facilitates seamless transitions between aerial and terrestrial modes, advancing novel motion and actuation strategies in multi-modal flapping-wing robotics.
@@
An SE(3)-based controller [20] ensures precise trajectory tracking and seamless mode transition. To validate the effectiveness of T3, we developed a fully functional prototype and conducted a series of real-world experiments, along with benchmark comparisons. The results demonstrate the exceptional performance of both the robot and its controller, highlighting the potential of multi-modal flapping-wing technologies for future aerial-ground robotic applications.
The contributions of the proposed robot T3 are summarized:
  • A novel bio-inspired elastic passive leg, modeled after the legs of juvenile water striders during their locomotion on the water’s surface, effectively transmits the vibrations of the flapping wings. This design leverages the vibrations generated during flapping-wing actuation to drive the elastic passive leg, enabling terrestrial locomotion for the first time.
  • A novel tailless triple-flapping-wing robot integrated with bio-inspired elastic passive legs, enabling both aerial flight and ground crawling, along with seamless mode transition. Remarkably, these capabilities are achieved without the need for additional actuators dedicated solely to terrestrial locomotion.
  • An SE(3)-based controller that achieves precise tracking of aerial trajectories without active yaw control, along with seamless mode transitions.
  • A series of real-world experiments, and benchmark comparisons that demonstrate the outstanding performance of both the robot and its controller.
II Design of Multi-model Robot
@@ II-A Bio-inspired Elastic Passive Leg
Water striders are exceptional insects known for their ability to generate propulsive vortices through rapid stroking motions of their long, specialized legs. This unique mechanism results in the formation of small whirlpools on the water’s surface, with the generated horseshoe vortex providing the force necessary to propel the insect forward. Exploiting this hydrodynamic effect, water striders are capable of gliding across the surface at impressive speeds [16] [17] [18], as depicted in Fig. 2(a), (c). A detailed structural analysis reveals that the insect possesses three pairs of elongated legs. Among these, the middle pair plays a crucial role in propulsion due to their larger size relative to the smaller front pair, which contributes minimally to the propulsion process. Additionally, the hind pair serves to stabilize the insect’s body, ensuring balance during locomotion and facilitating smooth gliding across the water’s surface [19], as depicted in Fig. 2(a).
In developing a terrestrial locomotion system, we propose a bio-inspired elastic passive leg mechanism that emulates the leg dynamics of water striders. The structural configuration is illustrated in Fig. 2(b). The system generates propulsive force through a motor-driven mechanism that actuates flapping-wing structures, while inducing vibrations. These vibrations are effectively transmitted through a lightweight carbon fiber rod structure to the support legs in contact with the ground. The support legs are designed with inherent elasticity to replicate the rapid stroking motion characteristic of water striders.
Both the primary leg structure and supporting limbs incorporate a degree of flexibility, optimizing energy transfer efficiency and enabling dynamic locomotion akin to the rapid leg movements of water striders on the water’s surface. By leveraging nature-inspired locomotion principles, this design achieves efficient propulsion while maintaining stability, offering insights into high-performance terrestrial robotic mobility.
A pair of symmetrically arranged bio-inspired elastic passive legs is implemented to replicate the primary propulsion mechanism exhibited by the middle legs of a water strider. These legs generate the majority of the propulsive force, emulating the rapid and efficient stroking motion that enables the insect to traverse the water surface. Additionally, a single centrally positioned bio-inspired elastic passive leg is integrated at the rear to enhance stability, analogous to the function of the hind legs in water striders, as illustrated in Fig. 2(a).
Unlike the water strider’s natural configuration, which employs two hind legs, this design utilizes a single stabilizing leg to reduce ground friction, improve maneuverability, and enhance overall walking stability. This simplification streamlines the control system and provides greater adaptability in movement, ultimately contributing to the efficiency and robustness of the locomotion system.
@@ II-B Fly Mode of Triple-flapping-wing Robot
Previous research (Delfly series [21] [22]) has achieved controlled flight in two-winged and four-winged flapping-wing vehicles, though most designs require either additional actuators to twist wing trailing edges or pre-tilted wing arrangements to generate yaw torque - both approaches that compromise lift efficiency. Notably, yaw motion does not affect the upright stability of flapping-wing vehicles [23], and the flapping mechanism itself does not inherently induce vehicle rotation, since flapping-wing motion does not generate reactive torque similar to rotating rotor blades when producing lift. These characteristics enable the possibility of autonomous flight in triple-flapping-wing system without active yaw control.
Figure 3: The block diagram of T3 control system in fly mode.
Stable closed-loop control of pitch and roll angles is the core of achieving hovering flight for tailless flapping-wing robots. This control is implemented through the cascaded structure of angular velocity PID controller and angle PID controller. Due to the lack of active yaw control, position control of this robot, in addition to position PID and velocity PID, also requires the allocation of velocity PID outputs based on the robot’s yaw angle. The flight control process of the robot under remote input is illustrated in Fig. 3.
@@
However, the inherent issue of yaw angle drift poses a significant challenge to position control systems that rely exclusively on on-board control algorithms, potentially leading to a progressive degradation of flight stability over extended operational durations. Furthermore, the implementation of cascade controllers for trajectory tracking tasks inevitably introduces complexities in parameter tuning processes and a heightened risk of compromised flight stability.
@@ II-C System Architecture and Components
Each flapping-wing module is rigidly connected to the main carbon rod frame of the robot via a $1.4\,\mathrm{mm}$ wide square carbon rod bracket. The electronic components of the robot body are illustrated in Fig. 6. The entire robot is powered by a $380\,\mathrm{mAh}$ 1S lithium battery and is equipped with a self-developed STM32 flight control board. This board integrates a MPU6500 gyroscope, a SPL06 barometer, a coreless motor driver circuit, and an optical flow communication circuit. Additionally, the board features interfaces for a Bluetooth module and a SBUS receiver, which are used for communication with the host computer and real-time signal reception from the remote control, respectively.
Figure 6: The block diagram of T3’s electronic system.
@@ III-A Crawling and Terrestrial Path Tracking
As shown in Fig. 7(a), a bottom limit bracket is used to isolate the vibrations of the flapping-wing modules. The terminal part of the limit bracket only constrains the tangential vibration $V_{T}$ at the bottom of the flapping-wing module while releasing the radial vibration $V_{R}$, ensuring that only radial vibrations affect the robot’s body during flapping motion. Additionally, the left and right arms of the chassis limit bracket are not in the same plane as those of the main fixed bracket of the flapping-wing module but are inclined at an angle $\beta$ toward the rear flapping-wing module. Moreover, the crawling leg of the rear flapping-wing module is higher $\varDelta H$ than the others. These two parameters amplify the driving effect of the front tangential vibration $V_{TF}$ while mitigating the impact of the rear tangential vibration $V_{TB}$.
After optimizing the chassis structure, increasing the motor output of the right flapping-wing module enables the robot to generate forward motion and counterclockwise rotation, which is illustrated in Fig. 7(b). Coordinating with the left flapping-wing module allows for differential control, achieving ground movements such as forward motion and turning. The robot has also successfully completed the ground trajectory tracking experiment. As shown in Fig. 4(a), the yaw angle of the ground crawling motion was sampled by the host computer, and the target yaw is given by Eq. (18). The synthesized image of the 8-shaped trajectory tracking experiment is illustrated in Fig. 8(a).
$\tan\psi_{T}=\frac{y_{T}-y}{x_{T}-x}$ (18)
Figure 8: (a) Ground tracking 8-shaped trajectory synthetic image, (b) Images of the self-righting process within 0.5 seconds.

### v2

# Tailless Flapping-Wing Robot with Bio-Inspired Elastic Passive Legs for Multi-Modal Locomotion

@@ Abstract
Flapping-wing robots offer significant versatility; however, achieving efficient multi-modal locomotion remains challenging. This paper presents the design, modeling, and experimentation of a novel tailless flapping-wing robot with three independently actuated pairs of wings. Inspired by the leg morphology of juvenile water striders, the robot incorporates bio-inspired elastic passive legs that convert flapping-induced vibrations into directional ground movement, enabling locomotion without additional actuators. This vibration-driven mechanism facilitates lightweight, mechanically simplified multi-modal mobility. An SE(3)-based controller coordinates flight and mode transitions with minimal actuation. To validate the robot’s feasibility, a functional prototype was developed, and experiments were conducted to evaluate its flight, ground locomotion, and mode-switching capabilities. Results show satisfactory performance under constrained actuation, highlighting the potential of multi-modal flapping-wing designs for future aerial-ground robotic applications. These findings provide a foundation for future studies on frequency-based terrestrial control and passive yaw stabilization in hybrid locomotion systems.
Index Terms:
Aerial Systems: Mechanics and Control, Biologically-Inspired Robots.
@@ I Introduction
The study of flapping-wing robots has garnered significant attention due to their promising applications in areas such as disaster rescue, environmental monitoring, urban exploration, and entertainment performances [1, 2, 3, 4]. However, these robots often face limitations in maneuverability and adaptability in complex environments [5, 6, 7]. To address these challenges, researchers have developed multi-modal flapping-wing robots capable of both aerial and terrestrial locomotion. This advancement enables robots to traverse diverse terrains and perform a wider range of tasks, thereby enhancing their versatility and effectiveness in real-world applications.
Figure 1: Multi-modal tailless triple-flapping-wing robot.
Integrating mechanisms from ground robots has become a common approach to achieving terrestrial mobility for multi-modal flapping-wing robots, due to their well-established reliability. However, equipping these robots with wheels [8, 9] and walking mechanisms [7, 9, 10, 11] often requires additional actuators compared to their flight-only configurations. This not only complicates the control system but also reduces the robot’s reliability and increases manufacturing and maintenance costs [12]. Moreover, this approach often results in inefficient multi-modal locomotion and switching.
Regarding aerial mobility, there is a growing trend in biomimetic research to achieve high levels of mimicry of natural organisms [5, 6]. For instance, some designs replicate insects by employing two sets of flapping mechanisms and two servos, while others mimic insects with two pairs of independent wings and two rotors [8]. Additionally, various bio-inspired designs have been developed, drawing inspiration from larger birds [10], hummingbirds [13], flying squirrels [14], ladybird beetles [15], and others. To achieve high levels of biomimicry and full attitude control in the air, existing studies often involve complex designs with numerous actuators, which can lead to inefficiency in multi-modal locomotion.
@@
Building upon previous analysis, we present a multi-modal tailless flapping-wing robot equipped with bio-inspired elastic passive legs modeled after those of juvenile water striders, which are used for locomotion on the water’s surface [16, 17, 18]. This design leverages the robot’s wing vibrations to enable terrestrial locomotion, integrating various functionalities such as vertical takeoff, multi-degree-of-freedom flight, self-righting, terrestrial locomotion, and seamless mode transitions, all achieved with just three actuators. The robot features three sets of annularly symmetric flapping-wing actuators for propulsion, with a total weight of $37.4\,\mathrm{g}$.A similar configuration utilizing three pairs of flapping wings for aerial mobility was previously demonstrated by de Wagter [19]. In flight mode, it reaches a maximum speed of $5.5\,\mathrm{m/s}$, with a flight endurance of $6.5\,\mathrm{minutes}$. In terrestrial mode, it achieves a top speed of $5.4\,\mathrm{cm/s}$ and can sustain controlled locomotion for up to $28\,\mathrm{minutes}$.
Figure 2: (a) Comparison between the juvenile water strider gliding on the water’s surface using active leg strokes [16] [17] [20], and the proposed robot using vibration-driven crawling, (b) Proposed bio-inspired elastic passive leg used in the robot, structurally inspired by the spatial arrangement of insect legs, (c) Side view of a water strider’s leg during locomotion [18].
@@
This design effectively addresses key challenges in both terrestrial and aerial mobility. By incorporating bio-inspired elastic passive legs, our approach utilizes wing vibrations for ground movement, eliminating the need for additional actuators. This not only simplifies the control system but also enhances reliability while reducing manufacturing and maintenance costs. In terms of aerial mobility, the three-winged configuration optimizes thrust generation and stability while maintaining a lightweight structure. Unlike conventional designs, which rely on multiple actuators for full-attitude control, our method reduces system complexity by using only three actuators to achieve control over roll, pitch, and thrust. While the vehicle does not provide active yaw control, it maintains sufficient maneuverability in the most critical degrees of freedom for stable and agile flight. This innovative approach improves the efficiency of multi-modal locomotion and facilitates seamless transitions between aerial and terrestrial modes, advancing novel motion and actuation strategies in multi-modal flapping-wing robotics.
@@
An SE(3)-based controller [21] ensures precise trajectory tracking and seamless mode transition. To validate the effectiveness of the robot, we developed a fully functional prototype and conducted a series of real-world experiments, along with benchmark comparisons. The results demonstrate the acceptable performance of both the robot and its controller, highlighting the potential of multi-modal flapping-wing technologies for future aerial-ground robotic applications.
The contributions of the proposed robot are summarized:
  • A novel vibration-driven terrestrial locomotion mechanism using bio-inspired elastic passive legs. The leg structure is inspired by the spatial configuration of juvenile water striders, but the locomotion principle is based on transmitting flapping-wing-induced vibrations through curved elastic legs achieving directional ground movement without additional actuators.
  • A tailless triple-flapping-wing robot integrating the above mechanism to enable both aerial flight and ground crawling. The system achieves seamless mode transitions between modalities, while maintaining low mechanical complexity and actuation count.
  • An SE(3)-based controller capable of stabilizing flight and executing trajectory tracking in underactuated conditions. Despite the absence of active yaw control, the system demonstrates acceptable tracking performance during experimental evaluations.
  • A set of real-world experiments validating the robot’s multi-modal capabilities. Benchmark comparisons and endurance tests illustrate the feasibility and effectiveness of the integrated design in both aerial and terrestrial domains.
II Design of Multi-modal Robot
@@ II-A Bio-inspired Elastic Passive Leg
Water striders are exceptional insects known for their ability to generate propulsive vortices through rapid stroking motions of their long, specialized legs. This unique mechanism results in the formation of small whirlpools on the water’s surface, with the generated horseshoe vortex providing the force necessary to propel the insect forward. Exploiting this hydrodynamic effect, water striders are capable of gliding across the surface at impressive speeds [16, 17, 18], as depicted in Fig. 2 (a) and (c). A visual inspection reveals that the insect possesses three pairs of elongated legs, among which the middle pair plays a crucial role in propulsion due to their relatively larger size. The hind legs mainly serve to stabilize the body and ensure balance during motion [20].
Inspired by the morphology and functional leg arrangement of water striders, we propose a bio-inspired elastic passive leg structure for terrestrial locomotion. The structural configuration is illustrated in Fig. 2 (b). In our design, propulsion is achieved not by large-amplitude active strokes as in water striders, but through small-amplitude, motor-induced flapping wing vibrations transmitted to the ground via asymmetric elastic legs. This vibration-driven mechanism, resembling that used in Kilobot [22] developed by Rubenstein et al., enables directional movement via energy transfer through lightweight carbon fiber rods and compliant support legs.
While the leg structure draws inspiration from the spatial layout and flexibility of the insect’s legs, the locomotion principle is based on vibration rather than fluid-dynamic propulsion. A pair of symmetrically arranged elastic legs replicates the structural role of the middle legs, serving as the primary contact and propulsion interface. In addition, a single passive leg positioned at the rear enhances stability, analogous in function to the hind legs of the water strider, but implemented as a simplification for reduced friction and greater control robustness.
By leveraging this hybrid bio-inspired design, structurally based on the water strider and mechanically aligned with vibration-driven motion, we achieve an efficient and stable terrestrial locomotion strategy suitable for lightweight underactuated robots.
Figure 3: The block diagram of the robot control system in fly mode.
@@ II-B Fly Mode of Triple-flapping-wing Robot
Previous research (Delfly series [23, 24]) has achieved controlled flight in two-winged and four-winged flapping-wing vehicles, though most designs require either additional actuators to twist wing trailing edges or pre-tilted wing arrangements to generate yaw torque - both approaches that compromise lift efficiency. Notably, yaw motion does not affect the upright stability of flapping-wing vehicles [25], and the flapping mechanism itself does not inherently induce continuous vehicle rotation, since flapping-wing mechanisms generate only transient inertial torques, in contrast to the sustained reactive torque produced by rotors during lift generation. These characteristics enable the possibility of autonomous flight in triple-flapping-wing system without active yaw control.
Stable closed-loop control of pitch and roll angles is the core of achieving hovering flight for tailless flapping-wing robots. This control is implemented through the cascaded structure of angular velocity PID controller and angle PID controller. Due to the lack of active yaw control, position control of this robot, in addition to position PID and velocity PID, also requires the allocation of velocity PID outputs based on the robot’s yaw angle. The flight control process of the robot under remote input is illustrated in Fig. 3. The onboard control loop runs at 100 Hz, and sensor data are sampled at 200 Hz. These rates are sufficient to support reasonably stable flight and consistent attitude responses during experimental testing.
@@
However, the inherent issue of yaw angle drift poses a significant challenge to position control systems that rely exclusively on onboard control algorithms, potentially leading to a progressive degradation of flight stability over extended operational durations. Furthermore, the implementation of cascade controllers for trajectory tracking tasks inevitably introduces complexities in parameter tuning processes and a heightened risk of compromised flight stability.
@@ II-C System Architecture and Components
Each flapping-wing module is rigidly connected to the main carbon rod frame of the robot via a $1.4\,\mathrm{mm}$ wide square carbon rod bracket. The electronic components of the robot body are illustrated in Fig. 6. The entire robot is powered by a $380\,\mathrm{mAh}$ 1S lithium battery and is equipped with a self-developed STM32 flight control board. This board integrates a MPU6500 gyroscope, a SPL06 barometer, a coreless motor driver circuit, and an optical flow communication circuit. The optical flow sensor provides velocity measurements relative to the ground, which are fused in the EKF to improve position and velocity estimation during low-altitude flight. Additionally, the board features interfaces for a Bluetooth module and a SBUS receiver, which are used for communication with the host computer and real-time signal reception from the remote control, respectively.
Figure 6: The block diagram of the robot’s electronic system.
@@ III-A Crawling and Terrestrial Path Tracking
To address these challenges, our design introduces deliberate structural asymmetries that bias the direction of motion. Specifically, the left and right arms of the chassis limit bracket are inclined backward by an angle $\beta$ relative to the main wing brackets, and the rear passive leg is positioned higher by a vertical offset $\varDelta H$ compared to the front legs. These features suppress the contribution of rear-leg-induced vibration while enhancing the forward propulsion generated by the front legs. As a result, the robot consistently exhibits forward crawling under symmetric flapping actuation, backward crawling is not possible under current design.
@@
The directionality of the motion is therefore determined primarily by mechanical design, rather than control-based modulation. While the locomotion system has not been explicitly optimized for frequency-dependent behavior, we have not observed any reversal or deviation in direction across the tested range of operating frequencies. This indicates that the asymmetry-driven structure offers reliable and robust bias for vibration-induced crawling under the current actuation scheme.
As shown in Fig. 7 (a), a bottom limit bracket with integrated sliding slots is used to isolate and direct the vibrations generated by the flapping-wing modules. In the final design, the sliding slots are oriented parallel to their attached support bars, which constrains the tangential vibration component $V_{T}$ while allowing the radial component $V_{R}$ to propagate toward the robot body. This configuration ensures that primarily radial vibrations contribute to body motion during flapping, thereby enhancing forward crawling performance.
In addition, the left and right arms of the chassis limit bracket are inclined at an angle $\beta$ toward the rear flapping-wing module, rather than being coplanar with the main fixed bracket. Furthermore, the rear crawling leg is elevated by a height difference $\varDelta H$ relative to the others. These two asymmetry parameters amplify the driving effect of the front tangential vibration $V_{TF}$ while mitigating the influence of the rear tangential vibration $V_{TB}$, resulting in more effective forward locomotion.
Finally, the legs themselves adopt a curved (curvy) shape in the final prototype. This curvature improves vibration transmission by introducing compliant deformation and enhancing contact dynamics with the ground. Compared to earlier flat-leg variants, the curvy design leads to more efficient vibration-driven propulsion.
Figure 8: (a) Ground tracking 8-shaped trajectory synthetic image, (b) Images of the self-righting process within $0.5\,\mathrm{seconds}$.
After optimizing the chassis structure, increasing the motor output of the right flapping-wing module enables the robot to generate forward motion and counterclockwise rotation, which is illustrated in Fig. 7 (b). Coordinating with the left flapping-wing module allows for differential control, achieving ground movements such as forward motion and turning. The robot has also completed the ground trajectory tracking experiment. As shown in Fig. 4 (a), the yaw angle of the ground crawling motion $\psi$ is sampled by the host computer, and the target yaw $\psi_{T}$ is given by Eq. (18), where $(x_{t},y_{t})$ refers to the target position, and $(x,y)$ shows the sampled position acquired from either optical flow or motion capture system. We propose a dual-layer PID controller based on the differential steering strategy, which consists of two modules: a linear distance PID and a planar yaw PID. The former generates the output $o_{d}$ based on the Euclidean distance error between the target position and the current pose $(x,y)$, while the latter produces the output $o_{\psi}$ by calculating the angular error between $\psi_{T}$ and $\psi$, and the actuation frequencies of the left and right flapping-wing modules are derived from $f_{1}=o_{d}+o_{\psi},f_{2}=o_{d}-o_{\psi}$. The synthesized image of the 8-shaped trajectory tracking experiment is illustrated in Fig. 8 (a).
$\tan\psi_{T}=\frac{y_{T}-y}{x_{T}-x}.$ (18)